OpenStudy (anonymous):

What is the common difference of a 43-term arithmetic sequence where the first term is -13 and the sum is 9,374?

Parth (parthkohli):

\[9374 = {43 \over 2}\left(-13 + n \right) \]Solve for \(n\) to get the last number in this sequence.

hartnn (hartnn):

use this equation \(S_n=n/2(2a+(n-1)d)\) n= 43,a=1st term=-13 Sn=9374

Parth (parthkohli):

\[d = {a_x - a_y \over x - y} \]

OpenStudy (anonymous):

ok what answers did u get? i think i got an asnwer

hartnn (hartnn):

11 ?

OpenStudy (anonymous):

that's what i got too. ok cool thanks for your help

hartnn (hartnn):

welcome :)